# local non-bossiness

### v3

@@ Abstract.
The student-optimal stable mechanism (DA), the most popular mechanism in school choice, is the only one that is both stable and strategy-proof. However, when DA is implemented, a student can change the schools of others without changing her own. We show that this drawback is limited: a student cannot change her classmates while remaining in the same school. We refer to this new property as local non-bossiness and use it to provide a new characterization of DA that does not rely on stability. Furthermore, we show that local non-bossiness plays a crucial role in providing incentives to be truthful when students have preferences over their colleagues. As long as students first consider the school to which they are assigned and then their classmates, DA induces the only stable and strategy-proof mechanism. There is limited room to expand this preference domain without compromising the existence of a stable and strategy-proof mechanism.
Keywords: School Choice - Preferences over Colleagues - Local Non-bossiness
@@
We are indebted to Keisuke Bando, Haluk Ergin, and William Thomson for their suggestions. We thank (in alphabetic order) Julien Combe, Gabriela Denis, Juan Dubra, Federico Echenique, Onur Kesten, SangMok Lee, Jordi Massó, Ran Shorrer, Camilo J. Sirguiado, and Bumin Yenmez for their valuable comments and suggestions, as well as the conference/seminar participants at 35th Stony Brook International Conference on Game Theory, LACEA-LAMES 2024, and Universidad de Montevideo. Financial support from ANII, FCE_1_2023_1_176072, is gratefully acknowledged.
E. Duque, London School of Economics and Political Science (e.g.duque@lse.ac.uk)
@@
J. P. Torres-Martínez, Department of Economics, University of Chile (juan.torres@fen.uchile.cl)
@@ 1. Introduction
In this paper, we introduce a new incentive property satisfied by ${\rm{DA}}$ that limits its bossiness. The new property, called local non-bossiness, says that a student cannot change her classmates without changing the school to which she is assigned. Thus, when a locally non-bossy mechanism is used, a student’s capacity to modify others’ assignment without changing her school is limited to those who are not assigned to the same school. This includes the case where the student is not assigned: she cannot modify the set of unassigned students while remaining without school. Local non-bossiness is therefore non-trivial even in a one-to-one matching problem. We use local non-bossiness, along with a set of axioms that do not include stability, to characterize the DA mechanism. We also study a model with externalities and show that local non-bossiness ensures the existence of a stable and strategy-proof mechanism when students care first about the school to which they are assigned, and then about their classmates.
Our analysis begins by showing that ${\rm{DA}}$ is locally non-bossy (Theorem 1) and that this property is independent of both stability or strategy-proofness. Furthermore, any locally non-bossy and strategy-proof mechanism is locally group strategy-proof, in the sense that no coalition of classmates can manipulate it to improve the situation for at least one of its members without hurting any other member. In particular, ${\rm{DA}}$ is the only stable and locally group strategy-proof mechanism (Corollary 1). We also study the relationships between local non-bossiness, local group strategy-proofness, and other incentive properties (the Figure 1 summarizes our results).
@@
However, even on this restricted domain, a stable and strategy-proof mechanism may not exist (Duque and Torres-Martínez, 2023). We further restrict the preference domain to the family of school-lexicographic preferences over colleagues, assuming that students care first about the school and then only about their classmates. That is, each student is indifferent among all the matchings in which she is assigned to the same school with the same classmates. There are no restrictions on the order in which she ranks two matchings that assign her to the same school but with different classmates. In this context, we demonstrate that applying ${\rm{DA}}$ to the school rankings underlying school-lexicographic preferences over colleagues induces a stable and strategy-proof mechanism (Theorem 3). Intuitively, a student may want to misreport her preferences to either change her school or maintain it and change her classmates. The first reason for misreporting preferences is already present in classical school choice problems, and avoiding it relates to ensuring strategy-proofness. The second one emerges in the presence of school-lexicographic preferences, and avoiding it relates to guaranteeing local non-bossiness. Therefore, the local non-bossiness and strategy-proofness of ${\rm{DA}}$ are key to ensuring the existence of a stable and strategy-proof mechanism in this context.
@@
Consequently, the incompatibility between stability and strategy-proofness in contexts where students prioritize their own school is not caused by the existence of preferences over the assignment of others per se, but by the fact that these preferences extend beyond their classmates. Evidently, it is reasonable to maintain the dependence of preferences on classmates because it is a well-documented empirical phenomenon in school choice (Rothstein, 2006; Abdulkadiroğlu et al., 2020; Allende, 2021; Che et al., 2022; Beuermann et al., 2023; Cox et al., 2023).
@@
Our results have practical implications for admission systems based on the ${\rm{DA}}$ mechanism. They show that ${\rm{DA}}$ still performs well when students care about the assignments of others, as long as they first consider their assigned school and then their classmates. More precisely, when DA is implemented and students are only required to report strict rankings of schools, it is a weakly dominant strategy for each student with school-lexicographic preferences over colleagues to report the school ranking induced by her true preferences. Furthermore, despite the variety of mechanisms that could be defined using information on preferences for schools and classmates, ${\rm{DA}}$ induces the only stable and strategy-proof mechanism in this domain (Corollary 3). It is important to note that there is limited room to relax the assumption of school-lexicographic preferences over colleagues. Indeed, it is enough for the preferences of just one student to be school-lexicographic but not school-lexicographic over colleagues to prevent the existence of a stable and strategy-proof mechanism (see Section 5).
@@
The concept of non-bossiness was first introduced by Satterthwaite and Sonnenschein (1981) and has been studied extensively in the context of the assignment of indivisible goods. [^4] We refer to the critical survey by Thomson (2016) for a detailed discussion of the multiple interpretations of non-bossiness and its implications. Non-bossiness plays an essential role in avoiding coalitional manipulability, as group strategy-proofness is equivalent to strategy-proofness and non-bossiness when preferences are strict (Papái, 2000). Moreover, for a given pair of schools’ priorities and capacities, Ergin (2002) shows that ${\rm{DA}}$ is non-bossy if and only if it is Pareto efficient. Kojima (2010) demonstrates the incompatibility between non-bossiness and stability in college admission problems, a scenario where both sides of the market may act strategically. However, when only one side of the market reports preferences, as in school choice problems, Afacan and Dur (2017) show that the school-optimal stable mechanism is non-bossy for the students. We contribute to this strand of the literature by highlighting that the bossiness of ${\rm{DA}}$ is limited: when it is implemented, no student can change her classmates without also changing her assigned school. We also show that this property implies that no group of students assigned to the same school can manipulate DA to either improve the assignment of at least one of them without hurting any other member (locally group strategy-proofness), or maintain the school and change the other classmates (locally group non-bossiness).
@@
Although Dutta and Massó (1997) demonstrate that the set of stable matchings when preferences are school-lexicographic is the same as without externalities, Duque and Torres-Martínez (2023) show that a stable and strategy-proof mechanism may not exist. We contribute to this strand of the literature by identifying a new preference domain where this incompatibility does not hold. In particular, when students first consider the school to which they are assigned and then only their classmates, ${\rm{DA}}$ induces the only stable and strategy-proof mechanism. Moreover, it is the local non-bossiness of ${\rm{DA}}$ that underlies this result.
The rest of the paper is organized as follows. Section 2 describes the classical school choice problem and the concepts of local non-bossiness and local group strategy-proofness. In Section 3 we show that ${\rm{DA}}$ is locally non-bossy. Section 4 provides an axiomatic characterization of DA. In Section 5 we introduce externalities in our model, and we show that ${\rm{DA}}$ induces a stable and strategy-proof mechanism in the domain of school-lexicographic preferences over colleagues. Section 6 contains the final remarks. Omitted proofs and examples are left to the appendices.
@@ 2. Model
When a mechanism is locally non-bossy, no student can modify her classmates without also changing the school to which she is assigned. The definition contemplates a situation where the student is unassigned. In this case, the student cannot modify the set of unassigned students without getting a seat in a school. This implies that local non-bossiness is non-trivial even in a one-to-one matching problem.
@@
When a mechanism is locally group strategy-proof, no coalition of classmates can manipulate it to improve the welfare of at least one of its members.
In Appendix A we analyze the relationships between local non-bossiness, local group strategy-proofness, and other incentive properties (the Figure 1 summarizes our results). Specifically, we demonstrate that local non-bossiness is independent of both stability and strategy-proofness, while local group strategy-proofness is weaker than group strategy-proofness and stronger than strategy-proofness. Furthermore, any locally non-bossy and strategy-proof mechanism is locally group strategy-proof (Lemma 1), but a locally group strategy-proof mechanism might violate local non-bossiness. This last result contrasts with the well-known equivalence between group strategy-proofness and the combination of non-bossiness and strategy-proofness (Papái, 2000). Additionally, we show in Lemma 2 that any locally non-bossy and strategy-proof mechanism satisfies a local notion of group non-bossiness (Afacan, 2012), and in Lemma 3 that local non-bossiness is independent of weak non-bossiness (Bando and Imammura, 2016). [^8] A notion related to non-bossiness is consistency (Ergin, 2002). A mechanism is consistent if after removing some students with their seats, the rest of the students remain in the same school when the mechanism is applied to the reduced market. It is easy to formulate a local notion of consistency and to show that a locally consistent mechanism is locally non-bossy.
@@ 3. Local non-bossiness of DA
However, when DA is implemented a student can modify the assignment of others without affecting her own (Ergin, 2002). Hence, one may wonder if there is a limit on the impact of these changes. In particular, under ${\rm{DA}}$, is it possible for a student to affect the set of students assigned to her school by reporting different preferences, without changing her school ? As the following result shows, the answer is negative.
@@
Proof.
@@
Fix a school choice context $[S,N,\succ,q]$ and w.l.o.g consider student $1$ and preference profiles $(P_{1},P_{-1}),(P^{\prime}_{1},P_{-1})\in{\mathcal{P}}$ such that $\overline{s}\equiv{\rm{DA}}_{1}(P_{1},P_{-1})={\rm{DA}}_{1}(P^{\prime}_{1},P_{-1})$. We will prove that ${\rm{DA}}_{\overline{s}}(P_{1},P_{-1})={\rm{DA}}_{\overline{s}}(P^{\prime}_{1},P_{-1})$. Let $\mu={\rm{DA}}(P_{1},P_{-1})$ and $\mu^{\prime}={\rm{DA}}(P^{\prime}_{1},P_{-1})$.
When $\overline{s}=s_{0}$, the result follows from Lemma 4 because no student who has not been assigned a seat in $\mu$ may be assigned a seat in $\mu^{\prime}$ (see Appendix B). When $\overline{s}\neq s_{0}$, suppose by contradiction that there exists $i_{1}\in\mu(\overline{s})$ such that $\mu^{\prime}(i_{1})\neq\overline{s}$. Without loss of generality, assume that $\mu^{\prime}(i_{1})P_{i_{1}}\overline{s}$ and denote by $I$ the non-empty set of students who prefer $\mu^{\prime}$ to $\mu$ [^9] When $\mu(\overline{s})\neq\mu^{\prime}(\overline{s})$, we have that $|\mu(\overline{s})|=|\mu^{\prime}(\overline{s})|=q_{\overline{s}}$ (see Lemma 4). Hence, $\mu(\overline{s})\setminus\mu^{\prime}(\overline{s})$ and $\mu^{\prime}(\overline{s})\setminus\mu(\overline{s})$ are non-empty sets. If we assume that $\overline{s}P_{i}\mu^{\prime}(i)$ for all $i\in\mu(\overline{s})\setminus\mu^{\prime}(\overline{s})$, then the stability of $\mu^{\prime}$ implies that $j\succ_{\overline{s}}i$ for all $j\in\mu^{\prime}(\overline{s})$ and $i\in\mu(\overline{s})\setminus\mu^{\prime}(\overline{s})$. Hence, the stability of $\mu$ ensures that $\mu(j)P_{j}\overline{s}$ for all $j\in\mu^{\prime}(\overline{s})\setminus\mu(\overline{s})$. Therefore, when there is no $i\in\mu(\overline{s})\setminus\mu^{\prime}(\overline{s})$ such that $\mu^{\prime}(i)P_{i}\overline{s}$, we can maintain our arguments by simply swapping the roles of $(P_{1},P_{-1})$ and $(P^{\prime}_{1},P_{-1})$.
@@
Let $G=(V,E)$ be the graph with nodes $V=I\,\cup\,\{j:\mu(j)=\mu^{\prime}(i)\,\,\mbox{for some}\,\,i\in I\}$ and directed edges $E=\{[i,j]:i\in I,\,\mu(j)=\mu^{\prime}(i)\}.$ Hence, the nodes of $G$ are the students who prefer $\mu^{\prime}$ to $\mu$, together with those who are assigned in $\mu$ to the schools to which students in $I$ are assigned in $\mu^{\prime}$. Moreover, there is an edge between every node $i\in I$ and all the students who are assigned to $\mu^{\prime}(i)$ in $\mu$.

### v4

@@ Abstract.
The student-optimal stable mechanism (DA), the most popular mechanism in school choice, is the only one that is stable and strategy-proof. However, when DA is implemented, a student can change the schools of others without changing her own. We show that this drawback is limited: a student cannot change her schoolmates while remaining in the same school. We refer to this new property as local non-bossiness and use it to provide a new characterization of DA that does not rely on stability. Furthermore, we show that local non-bossiness plays a crucial role in providing incentives to be truthful when students have preferences over their colleagues. As long as students first consider the school to which they are assigned and then their schoolmates, DA induces the only stable and strategy-proof mechanism. There is limited room to expand this preference domain without compromising the existence of a stable and strategy-proof mechanism.
Keywords: School Choice - Local Non-bossiness - Student-optimal stable mechanism - Preferences over Colleagues
@@
We are indebted to Keisuke Bando, Haluk Ergin, and William Thomson for their suggestions. We thank (in alphabetic order) Julien Combe, Gabriela Denis, Juan Dubra, Federico Echenique, Onur Kesten, SangMok Lee, Jordi Massó, Ran Shorrer, Camilo J. Sirguiado, Can Urgun and Bumin Yenmez for their valuable comments and suggestions, as well as the conference/seminar participants at 35th Stony Brook International Conference on Game Theory, LACEA-LAMES 2024, Universidad de Montevideo, London School of Economics, Universidad de Salamanca, and SEU meeting 2024. Financial support from ANII, FCE_1_2023_1_176072, is gratefully acknowledged.
E. Duque-Rosas, London School of Economics and Political Science (e.g.duque@lse.ac.uk)
@@
J. P. Torres-Martínez, Department of Economics, Faculty of Economics and Business, University of Chile (juan.torres@fen.uchile.cl)
@@ 1. Introduction
In this paper, we introduce a new incentive property satisfied by ${\rm{DA}}$ that limits its bossiness. The new property, called local non-bossiness, says that a student cannot change her schoolmates without changing the school to which she is assigned. Thus, when a locally non-bossy mechanism is used, a student’s capacity to modify others’ assignment without changing her school is limited to those who are not assigned to the same school. This includes the case where the student is not assigned: she cannot modify the set of unassigned students while remaining without school. Local non-bossiness is therefore non-trivial even in a one-to-one matching problem. We use local non-bossiness, along with a set of axioms that do not include stability, to characterize the DA mechanism. We also study a model with externalities and show that local non-bossiness ensures the existence of a stable and strategy-proof mechanism when students care first about the school to which they are assigned and then about their schoolmates.
Our analysis begins by showing that ${\rm{DA}}$ is locally non-bossy (Theorem 1) and that this property is independent of both stability or strategy-proofness. Furthermore, any locally non-bossy and strategy-proof mechanism is locally group strategy-proof, in the sense that no coalition of schoolmates can manipulate it to improve the situation of at least one of its members without hurting any other member. In particular, ${\rm{DA}}$ is the only stable and locally group strategy-proof mechanism (Corollary 1). We also study the relationships between local non-bossiness, local group strategy-proofness, and other incentive properties (the Figure 1 summarizes our results).
@@
However, even on this restricted domain, a stable and strategy-proof mechanism may not exist (Duque and Torres-Martínez, 2023). We further restrict the preference domain to the family of school-lexicographic preferences over colleagues, assuming that students care first about the school and then only about their schoolmates. That is, each student is indifferent among all the matchings in which she is assigned to the same school with the same colleagues. There are no restrictions on the order in which she ranks two matchings that assign her to the same school but with different schoolmates. In this context, we demonstrate that applying ${\rm{DA}}$ to the school rankings underlying school-lexicographic preferences over colleagues induces a stable and strategy-proof mechanism (Theorem 3). Intuitively, a student may want to misreport her preferences either to change her school or to maintain it and change her schoolmates. The first reason for misreporting preferences is already present in classical school choice problems, and avoiding it relates to ensuring strategy-proofness. The second one emerges in the presence of school-lexicographic preferences, and avoiding it relates to guaranteeing local non-bossiness. Therefore, the local non-bossiness and strategy-proofness of ${\rm{DA}}$ are key to ensuring the existence of a stable and strategy-proof mechanism in this context.
@@
Consequently, the incompatibility between stability and strategy-proofness in contexts where students prioritize their own school is not caused by the existence of preferences over the assignment of others per se, but by the fact that these preferences extend beyond their schoolmates. Evidently, it is reasonable to maintain the dependence of preferences on schoolmates because it is a well-documented empirical phenomenon in school choice (Rothstein, 2006; Abdulkadiroğlu et al., 2020; Allende, 2021; Che et al., 2022; Beuermann et al., 2023; Cox et al., 2023).
@@
Our results have practical implications for admission systems based on the ${\rm{DA}}$ mechanism. They show that ${\rm{DA}}$ still performs well when students care about the assignments of others, as long as they first consider their assigned school and then their colleagues. More precisely, when DA is implemented and students are only required to report strict rankings of schools, it is a weakly dominant strategy for each student with school-lexicographic preferences over colleagues to report the school ranking induced by her true preferences. Furthermore, despite the variety of mechanisms that could be defined using information on preferences for schools and schoolmates, ${\rm{DA}}$ induces the only stable and strategy-proof mechanism in this domain (Corollary 3). It is important to note that there is limited room to relax the assumption of school-lexicographic preferences over colleagues. Indeed, it is enough for the preferences of just one student to be school-lexicographic but not school-lexicographic over colleagues to prevent the existence of a stable and strategy-proof mechanism (see Section 5).
@@
The concept of non-bossiness was first introduced by Satterthwaite and Sonnenschein (1981) and has been studied extensively in the context of the assignment of indivisible goods. [^4] We refer to the critical survey by Thomson (2016) for a detailed discussion of the multiple interpretations of non-bossiness and its implications. Non-bossiness plays an essential role in avoiding coalitional manipulability, as group strategy-proofness is equivalent to strategy-proofness and non-bossiness when preferences are strict (Papái, 2000). Moreover, for a given pair of schools’ priorities and capacities, Ergin (2002) shows that ${\rm{DA}}$ is non-bossy if and only if it is Pareto efficient. Kojima (2010) demonstrates the incompatibility between non-bossiness and stability in college admission problems, a scenario where both sides of the market may act strategically. However, when only one side of the market reports preferences, as in school choice problems, Afacan and Dur (2017) show that the school-optimal stable mechanism is non-bossy for the students. We contribute to this strand of the literature by highlighting that the bossiness of ${\rm{DA}}$ is limited: when it is implemented, no student can change her schoolmates without also changing her assigned school. We also show that this property implies that no group of students assigned to the same school can manipulate DA to either improve the assignment of at least one of them without hurting any other member (locally group strategy-proofness), or maintain the school and change the other schoolmates (locally group non-bossiness).
@@
Although Dutta and Massó (1997) demonstrate that the set of stable matchings when preferences are school-lexicographic is the same as without externalities, Duque and Torres-Martínez (2023) show that a stable and strategy-proof mechanism may not exist. We contribute to this strand of the literature by identifying a new preference domain where this incompatibility does not hold. In particular, when students first consider the school to which they are assigned and then only their colleagues, ${\rm{DA}}$ induces the only stable and strategy-proof mechanism. Moreover, it is the local non-bossiness of ${\rm{DA}}$ that underlies this result.
The remainder of the paper is organized as follows. Section 2 describes the classical school choice problem and the concepts of local non-bossiness and local group strategy-proofness. In Section 3 we show that ${\rm{DA}}$ is locally non-bossy. Section 4 provides an axiomatic characterization of DA. In Section 5 we introduce externalities in our model, and we show that ${\rm{DA}}$ induces a stable and strategy-proof mechanism in the domain of school-lexicographic preferences over colleagues. Section 6 contains the final remarks. The omitted proofs and examples are left to the appendices.
@@ 2. Model
When a mechanism is locally non-bossy, no student can modify her schoolmates without also changing the school to which she is assigned. The definition contemplates a situation where the student is unassigned. In this case, the student cannot modify the set of unassigned students without getting a seat in a school. This implies that local non-bossiness is non-trivial even in a one-to-one matching problem.
@@
When a mechanism is locally group strategy-proof, no coalition of schoolmates can manipulate it to improve the welfare of at least one of its members.
In Appendix A we analyze the relationships between local non-bossiness, local group strategy-proofness, and other incentive properties (the Figure 1 summarizes our results). Specifically, we demonstrate that local non-bossiness is independent of both stability and strategy-proofness, while local group strategy-proofness is weaker than group strategy-proofness and stronger than strategy-proofness. [^8] Even well-known stable mechanisms do not satisfy local non-bossiness. This is the case of the median stable mechanism introduced by Klaus and Klijn (2006); see Example 7. Furthermore, any locally non-bossy and strategy-proof mechanism is locally group strategy-proof (Lemma 1), but a locally group strategy-proof mechanism might violate local non-bossiness. This last result contrasts with the well-known equivalence between group strategy-proofness and the combination of non-bossiness and strategy-proofness (Papái, 2000). Additionally, we show in Lemma 2 that any locally non-bossy and strategy-proof mechanism satisfies a local notion of group non-bossiness (Afacan, 2012), and in Lemma 3 that local non-bossiness is independent of weak non-bossiness (Bando and Imammura, 2016). [^9] A notion related to non-bossiness is consistency (Ergin, 2002). A mechanism is consistent if after removing some students with their seats, the rest of the students remain in the same school when the mechanism is applied to the reduced market. It is easy to formulate a local notion of consistency and to show that a locally consistent mechanism is locally non-bossy.
@@ 3. Local non-bossiness of DA
However, when DA is implemented a student can modify the assignment of others without affecting her own (Ergin, 2002). Hence, one may wonder if there is a limit on the impact of these changes. In particular, under ${\rm{DA}}$, is it possible for a student to affect the set of students assigned to her school by reporting different preferences, without changing her school ? As the following theorem shows, the answer is negative.
@@
To get an intuitive idea of the proof of Theorem 1, let $\mu={\rm{DA}}(P)$ and $\mu^{\prime}={\rm{DA}}(P^{\prime}_{i},P_{-i})$ be distinct matchings that give the same assignment to student $i$, namely $\overline{s}\equiv\mu(i)=\mu^{\prime}(i)$.
By appealing to the Rural Hospital Theorem, we can show that students who receive different assignments in $\mu$ and $\mu^{\prime}$ are in schools that fill their quotas in both matchings. Hence, when we compare $\mu^{\prime}$ to $\mu$, each student who changes assignments displaces another. These movements induce cycles of students $(i_{1},\ldots,i_{r})$, where each $i_{k}$ receives in $\mu^{\prime}$ the school that $i_{k+1}$ had in $\mu$ (modulo $r$). If there exists a student who prefers $\mu^{\prime}$ to $\mu$, [^10] If this does not occur, we can maintain our arguments by simply swapping the roles of $P$ and $(P^{\prime}_{i},P_{-i})$. the stability of $\mu$ and $\mu^{\prime}$ ensures that at least one of these cycles is a $\mu$-improving cycle—that is, a cycle $(i_{1},\ldots,i_{r})$ in which every $i_{k}$ prefers $\mu(i_{k+1})$ to $\mu(i_{k})$ (modulo $r$).
We define a graph $G$ by letting each student $j$ who prefers $\mu^{\prime}$ to $\mu$ point to those who are assigned in $\mu$ to the school $\mu^{\prime}(j)$. Previous arguments ensure that $G$ has a $\mu$-improving cycle. Since $\mu$ is the student-optimal stable matching under $P$, any $\mu$-improving cycle necessarily induces justified envy. By allowing students in $G$ to block the improving cycles contained in it, we construct a new $\mu$-improving cycle $(i^{*}_{1},\ldots,i^{*}_{r})$ in which justified envy is limited to students outside $G$—that is, to those who are assigned in $\mu$ to schools that, in $\mu^{\prime}$, do not receive anyone who prefers $\mu^{\prime}$ to $\mu$ [^11] The $\mu$-improving cycle $(i^{*}_{1},\ldots,i^{*}_{r})$ is not necessarily present in $G$. Moreover, since students in $N\setminus\{i\}$ do not change their preferences, we can show that only $i$ blocks $(i^{*}_{1},\ldots,i^{*}_{r})$. Hence, the students in $\mu^{\prime}(\overline{s})$ weakly prefer $\mu$ to $\mu^{\prime}$. This result implies that either $\mu^{\prime}(\overline{s})\subseteq\mu(\overline{s})$ or $\mu(\overline{s})\subseteq\mu^{\prime}(\overline{s})$. Finally, to conclude that $\mu(\overline{s})=\mu^{\prime}(\overline{s})$, we invoke the Rural Hospital Theorem to ensure that any school with different entering classes in $\mu$ and $\mu^{\prime}$ always fills its quota.
Proof of Theorem 1.
@@
Fix a school choice context $[S,N,\succ,q]$ and w.l.o.g consider student $1$ and preference profiles $(P_{1},P_{-1}),(P^{\prime}_{1},P_{-1})\in{\mathcal{P}}$ such that $\overline{s}\equiv{\rm{DA}}_{1}(P_{1},P_{-1})={\rm{DA}}_{1}(P^{\prime}_{1},P_{-1})$ and ${\rm{DA}}(P_{1},P_{-1})\neq{\rm{DA}}(P^{\prime}_{1},P_{-1})$. We will prove that ${\rm{DA}}_{\overline{s}}(P_{1},P_{-1})={\rm{DA}}_{\overline{s}}(P^{\prime}_{1},P_{-1})$.
When $\overline{s}=s_{0}$, the result follows from the fact that no student who is not assigned a seat in ${\rm{DA}}(P_{1},P_{-1})$ can be assigned a seat in ${\rm{DA}}(P^{\prime}_{1},P_{-1})$ (see Lemma 4 in Appendix B).
Assume that $\overline{s}\neq s_{0}$. Let $\mu={\rm{DA}}(P_{1},P_{-1})$ and $\mu^{\prime}={\rm{DA}}(P^{\prime}_{1},P_{-1})$. Notice that, without loss of generality, we can assume that the set $I$ of students who prefer $\mu^{\prime}$ to $\mu$ is non-empty. [^12] Since $\mu(1)=\mu^{\prime}(1)$ and $\mu\neq\mu^{\prime}$, when $I$ is an empty set there is at least one student who prefers $\mu$ to $\mu^{\prime}$. Hence, our arguments hold by swapping the roles of $(P_{1},P_{-1})$ and $(P^{\prime}_{1},P_{-1})$.
@@
Let $G=(V,E)$ be the graph with nodes $V=I\,\cup\,\{j:\mu(j)=\mu^{\prime}(i)\,\,\mbox{for some}\,\,i\in I\}$ and directed edges $E=\{[i,j]:i\in I,\,\mu(j)=\mu^{\prime}(i)\}.$ Hence, the nodes of $G$ are the students who prefer $\mu^{\prime}$ to $\mu$, together with those who are assigned in $\mu$ to the schools to which students in $I$ are assigned in $\mu^{\prime}$. Also, there is an edge between every $i\in I$ and all the students who are assigned to $\mu^{\prime}(i)$ in $\mu$.